# Experimental study of the distributions of off-diagonal scattering-matrix elements of quantum graphs with symplectic symmetry

### v1

@@ IV Distributions of the $S$-matrix elements
In the top panel of Fig. 4 the distributions of the $S$-matrix elements $S_{12}$ and $S_{1\bar{2}}$ are compared. They are obtained from measured transmission spectra from $P_{2}$ to $P_{1}$, denoted by $S_{12}$, and from $P_{\bar{2}}$ to $P_{1}$ denoted by $S_{1\bar{2}}$, respectively. In the middle and bottom panel these distributions are compared to those obtained for the numerically computed $S$ matrix. Further results are shown in Fig. 5, also for the distribution of $S_{11}$ for another size of absorption. The agreement is very good. Hence, we may conclude that the experiments indeed may be used to model the corresponding GSE graph, implying that we may use numerical computations as ancillary data for a better understanding of experimental ones. Indeed, the experimental results were at first a surprise, because the distributions for $S_{12}$ and $S_{1\bar{2}}$ clearly deviate from each other, in contradiction to RMT predictions with $V$ in Eq. (1) being a full matrix. The distributions are broader for scattering within a block, $S_{12}$, and sharper peaked around zero for $S_{1\bar{2}}$, implying that scattering between the spin-up and spin-down system is less likely than spin-preserving scattering. This observation led us to the conclusion that the discrepancies originate from the fact that in the GSE graph and microwave network only two of the nine vertices in one GUE graph are coupled to the corresponding two in the other GUE graph. To corroborate this, we also computed the $S$ matrix of the GUE graphs in which the subgraphs are coupled by four bonds. Results are shown in Figs. 6-7. For the graph with four coupling bonds the distributions of $S_{12}$ and $S_{1\bar{2}}$ agree well as predicted by RMT. Generally, for $a\neq b$ the distributions for the real and imaginary parts lie on top of each other for both cases, however the distributions for the case with two coupling bonds clearly deviate from those for four coupling bonds. On the contrary, for $a=b$ the distributions for the real and imaginary parts differ for each case, whereas the distributions of the two cases lie on top of each other. Accordingly, agreement with RMT predictions is found for the distributions of $S_{aa}$ [60], whereas deviations are observed for $S_{ab}\,(a\neq b)$ for the case with two coupling bonds. Note that the matrix elements $S_{a\bar{a}}$ vanish [43, 44].
@@ V Random-matrix simulations
For $c=1,\bar{1},2,\bar{2}$ the transmission coefficients are determined from measurements of the reflection spectra $S_{aa}$ via the relation $T_{a}=1-|\overline{S_{aa}}|^{2}$ [39] with $\overline{X}$ denoting the spectral average of $X$. The transmission coefficients associated with the fictitious channels are all set to $T_{f}$. Here we chose $\Lambda=25$ in the random-matrix simulations. These are related to absorption strength $\tau_{\rm abs}$ through the Weisskopf relation [122, 9], $\tau_{\rm abs}=2\Lambda T_{f}$. This parameter is used as fit parameter and is determined by comparing the RMT results for the distributions of the $S$-matrix elements, to the experimental ones. The resulting distributions for $S_{12}$ and $S_{1\bar{2}}$ are equal. In Fig. 8 the RMT results for the distributions of the modulus of $S_{11}$ (top) and of $S_{12}$ (bottom) are compared to those obtained from experiments and numerical computations. In the experiments absorption was varied as indicated in the legends of the panels by adding attenuators to corresponding bonds of the subgraphs. The agreement is quite good for that case, whereas the RMT curves clearly deviate from those for the microwave network and quantum graph for $S_{1\bar{2}}$, as can be inferred from the middle part of  Fig. 7. Good agreement is only found for the case with four coupling cables as demonstrated for the modulus of $S_{12}$ and $S_{1\bar{2}}$, respectively, in the top panel of Fig. 9.
We suspect that the discrepancies between the distributions for $S_{12}$ and $S_{1\bar{2}}$ observed in the experiments and the random-matrix simulations originate from the fact that in the RMT model a Hamiltonian of the form Eq. (1) is used in which $V$ is a full matrix. This corresponds to a coupling between all matrix elements of $H_{0}$ and $H_{0}^{\ast}$. However, in the microwave network and quantum graph, only two vertices of one subgraph are connected to the analogue ones in the other subgraph. Accordingly, we modified the Hamiltonian Eq. (1) entering the $S$ matrix, and set all entries of the matrix $V$ to zero except the elements with indices $(i_{0},\bar{j}),(\bar{i}_{0},j),(\bar{j},i_{0}),(j,\bar{i}_{0})$ for 5 values of $i_{0}$ and $j=1,\dots,N$ with $N=200$ in the $2N\times 2N$-dimensional Hamilton matrix $H$Eq. (1), that is, drastically reduced its rank. As demonstrated in the lower part of Fig. 9 then the agreement between the results from the RMT simulations for an ensemble of 300 $S$ matrices with $\tau_{\rm{abs}}=25$, the experiments, and the quantum graph numerics is good.
@@ VI Conclusions
We performed experiments with open microwave networks with symplectic symmetry, and analyzed distributions of the off-diagonal matrix elements $S_{ab}$ (which is equivalent to $S_{\bar{a}\bar{b}}$) and $S_{a\bar{b}}$ (which is equivalent to $S_{\bar{a}b}$) with $a\neq b$. We find that their distributions differ, in contradiction to RMT predictions. The discrepancies are attributed to the fact that the two subgraphs, that form the GSE graphs, are coupled by just one pair of coaxial cables. This is confirmed with numerical computations for GSE graphs, in which the subgraphs are coupled by 2 or more pairs of cables. Furthermore, we performed RMT simulations with a Hamiltonian of the form Eq. (1), where the matrix $V$ has a rank smaller than its dimension, corroborating our conjecture. Our findings imply that strictly speaking, because of the sparsity of the matrix $V$, the microwave networks and quantum graphs employed in the present work are no GSE graphs, even though their spectral properties agree well with those of random matrices from the GSE. Currently, we derive analytical expressions for the distributions of the off-diagonal $S$ matrix elements of quantum-chaotic scattering systems with symplectic symmetry composed of fully connected subsystems $H_{0}$ and $H_{0}^{\ast}$, which already is quite cumbersome. The next step will be to modify the model such that the analytical derivation using a similar approach becomes feasible for the case of partially connected subsystems, which is challenging.

### v2

@@ IV Distributions of the $S$-matrix elements
In the top panel of Fig. 4 the distributions of the $S$-matrix elements $S_{12}$ and $S_{1\bar{2}}$ are compared. They are obtained from measured transmission spectra from $P_{2}$ to $P_{1}$, denoted by $S_{12}$, and from $P_{\bar{2}}$ to $P_{1}$ denoted by $S_{1\bar{2}}$, respectively. In the middle and bottom panel these distributions are compared to those obtained for the numerically computed $S$ matrix. Further results are shown in Fig. 5, also for the distribution of $S_{11}$ for another size of absorption. The agreement is very good. Hence, we may conclude that the experiments indeed may be used to model the corresponding symplectic graph, implying that we may use numerical computations as ancillary data for a better understanding of experimental ones. Indeed, the experimental results were at first a surprise, because contrary to the property, that the distributions of all off-diagonal elements of the $S$ matrices of typical quantum-chaotic scattering systems with orthogonal or unitary symmetry – like GOE and GUE graphs –, respectively, are the same and coincide with RMT predictions [39, 74, 45, 99, 114, 66, 115], those for $S_{12}$ and $S_{1\bar{2}}$ clearly deviate from each other. The distributions are broader for scattering within a block, $S_{12}$, and sharper peaked around zero for $S_{1\bar{2}}$, implying that scattering between the spin-up and spin-down system is less likely than spin-preserving scattering. This observation led us to the conclusion that the discrepancies originate from the fact that in the symplectic graph and microwave network only two of the nine vertices in one GUE graph are coupled to the corresponding two in the other GUE graph. To corroborate this, we also computed the $S$ matrix of the symplectic graph in which the subgraphs are coupled by four bonds, whose spectral properties are exhibited in Fig. 3. Results for the distributions of the $S$-matrix elements are shown in Figs. 6-7. For the graph with four coupling bonds the distributions of $S_{12}$ and $S_{1\bar{2}}$ agree well. Generally, for $a\neq b$ the distributions for the real and imaginary parts lie on top of each other for both cases, however the distributions for the case with two coupling bonds clearly deviate from those for four coupling bonds. On the contrary, for $a=b$ the distributions for the real and imaginary parts differ for each case, whereas the distributions of the two cases lie on top of each other.
We also compared the experimental and numerical results for the $S$-matrix distributions with RMT predictions derived on the basis of the Heidelberg approach employing Hermitean random matrices from the GSE. As outlined in the following section, Sec. V, agreement with RMT predictions is only found for the distributions of $S_{aa}$ [60], whereas deviations are observed for $S_{ab}\,(a\neq b)$ for the case with two coupling bonds. Note that the matrix elements $S_{a\bar{a}}$ vanish [43, 44].
@@ V Random-matrix simulations
For $c=1,\bar{1},2,\bar{2}$ the transmission coefficients are determined from measurements of the reflection spectra $S_{aa}$ via the relation $T_{a}=1-|\overline{S_{aa}}|^{2}$ [39] with $\overline{X}$ denoting the spectral average of $X$. The transmission coefficients associated with the fictitious channels are all set to $T_{f}$. Here we chose $\Lambda=25$ in the random-matrix simulations. These are related to the absorption strength $\tau_{\rm abs}$ through the Weisskopf relation [124, 9], $\tau_{\rm abs}=2\Lambda T_{f}$. This parameter is used as fit parameter and is determined by comparing the RMT results for the distributions of the $S$-matrix elements, to the experimental ones. The resulting distributions for $S_{12}$ and $S_{1\bar{2}}$ are equal. In Fig. 8 the RMT results for the distributions of the modulus of $S_{11}$ (top) and of $S_{12}$ (bottom) are compared to those obtained from experiments and numerical computations. In the experiments absorption was varied as indicated in the legends of the panels by adding attenuators to corresponding bonds of the subgraphs. The agreement is quite good for that case, whereas the RMT curves clearly deviate from those for the microwave network and quantum graph for $S_{1\bar{2}}$, as can be inferred from the middle part of  Fig. 7. Good agreement is only found for the case with four coupling cables as demonstrated for the modulus of $S_{12}$ and $S_{1\bar{2}}$, respectively, in the top panel of Fig. 9.
As mentioned above, we suspect that the discrepancies between the distributions for $S_{12}$ and $S_{1\bar{2}}$ observed in the experiments and the random-matrix simulations based on the scattering matrix formalism Eq. (9) originate from the fact that in the RMT model Hermitean matrices from the GSE of the form Eq. (1) with a full $V$ matrix are used. This corresponds to a coupling between all matrix elements of $H_{0}$ and $H_{0}^{\ast}$. However, in the microwave network and quantum graph, only two vertices of one subgraph are connected to the analogue ones in the other subgraph. Accordingly, we modified the Hamiltonian Eq. (1) entering the $S$ matrix, and set all entries of the matrix $V$ to zero except the elements with indices $(i_{0},\bar{j}),(\bar{i}_{0},j),(\bar{j},i_{0}),(j,\bar{i}_{0})$ for 5 values of $i_{0}$ and $j=1,\dots,N$ with $N=200$ in the $2N\times 2N$-dimensional Hamilton matrix $H$, Eq. (1), that is, drastically reduced its rank. As demonstrated in the lower part of Fig. 9 then the agreement between the results from the RMT simulations for an ensemble of 300 $S$ matrices with $\tau_{\rm{abs}}=25$, the experimental ones, and the numerical ones is good.
@@ VI Conclusions
We performed experiments with open microwave networks with symplectic symmetry, and analyzed distributions of the off-diagonal matrix elements $S_{ab}$ (which is equivalent to $S_{\bar{a}\bar{b}}$) and $S_{a\bar{b}}$ (which is equivalent to $S_{\bar{a}b}$) with $a\neq b$. We find that their distributions differ, in contradiction to RMT predictions. The discrepancies are attributed to the fact that the two subgraphs, that form the symplectic graphs, are coupled by just one pair of coaxial cables. This is confirmed with numerical computations for symplectic graphs, in which the subgraphs are coupled by 2 or more pairs of cables. Furthermore, we performed RMT simulations with a Hamiltonian of the form Eq. (1), where the matrix $V$ has a rank smaller than its dimension, corroborating our conjecture. Our findings imply that strictly speaking, because of the sparsity of the matrix $V$, the microwave networks employed in the present work do not mimick quantum-chaotic $S$ matrices described by the scattering formalism Eq. (9) with $H$ a random matrix Eq. (1) from the GSE, even though their spectral properties agree well with GSE statistics. Yet, the properties of the $S$ matrix are well described by a random-matrix model of the form Eq. (9) with $V$ in Eq. (1) a matrix of rank lower than its dimension. Thus, the $S$ matrices of such microwave networks and of the corresponding quantum graphs exhibit universal properties. Currently, we derive analytical expressions for the distributions of the off-diagonal $S$-matrix elements of random matrices belonging to the symplectic universality class, composed of fully connected subsystems $H_{0}$ and $H_{0}^{\ast}$, i.e., random matrices from the GSE, which already is quite cumbersome. The next step will be to modify the model such that the analytical derivation using a similar approach becomes feasible for the case of partially connected subsystems, which is challenging.